# The Influence of Exclusion Zones on the Coexistence of Predator and Prey with an Allee Effect

### v1

@@ 5. Thin-limit analysis in 1D
We consider the following ansatze for the asymptotic expansions of $u_{1}$, $u_{2}$, and $v$ in small $a$:
@@
Note that the expansion for $v$ leads in order $a^{-1}$ since we expect the predator density to blow up as $a\to 0$ while the total predator population approaches a positive constant due to Theorem 5.1. Substituting (51) into (50) and comparing terms of like order in $a$, we immediately conclude that $v_{-1}$, $v_{0}$, and $u_{1,0}$ are all constant, and due to Theorem 5.1, $v_{-1}>0$. In order 1-1, we have $v_{1}^{\prime\prime}+(\gamma-\alpha u_{1,0})v_{-1}\,dx)\,dx=0$ with $v_{1}^{\prime}(0)=v_{1}^{\prime}(1)=0$, so we must have $u_{1,0}=\gamma/\alpha$.
Since $u_{1,0}=\gamma/\alpha$, we see that $u_{1,1}$ solves
$\begin{cases}d_{u}u_{1,1}^{\prime\prime}=\beta u_{1,0}v_{-1}=\frac{\beta\gamma}{\alpha}v_{-1}&0<x<1\\ u_{1,1}^{\prime}(0)=0&\\ u_{1,1}^{\prime}(1)=\frac{1}{L}u_{2,0}^{\prime}(0)\end{cases}$ (52)
The value of $v_{-1}$ is determined by the solvability condition:
$\frac{1}{L}u_{2,0}^{\prime}(0)=u_{1,1}^{\prime}(1)=\int_{0}^{1}u_{1,1}^{\prime\prime}\,dx=\frac{\beta\gamma}{d_{u}\alpha}v_{-1},$ (53)
leading to
$v_{-1}=\frac{d_{u}\alpha}{L\beta\gamma}u_{2,0}^{\prime}(0).$ (54)
The value $u_{2,0}^{\prime}(0)$ may be computed by solving the equation
$\begin{cases}\frac{d_{u}}{L^{2}}u_{2,0}^{\prime\prime}+f(u_{2,0})=0&0<x<1\\ u_{2,0}^{\prime}(1)=0&\\ u_{2,0}(0)=u_{1,0}(1)=\frac{\gamma}{\alpha}.\end{cases}$ (55)
With the value of $v_{-1}$ in hand, we can determine $u_{1,1}$ up to an additive constant:
$u_{1,1}(x)=\frac{1}{2L}u_{2,0}^{\prime}(0)x^{2}+c.$ (56)
@@
$\begin{cases}d_{v}v_{2}^{\prime\prime}=-\alpha u_{1,1}v_{-1}&0<x<1\\ v_{2}^{\prime}(0)=v_{2}^{\prime}(1)=0.\end{cases}$ (57)
We must have $\int_{0}^{1}v_{2}^{\prime\prime}\,dx=-\alpha v_{-1}\int_{0}^{1}u_{1,1}\,dx=0$. Thus, we find
$u_{1,1}(x)=\frac{1}{2L}u_{2,0}^{\prime}(0)\left(x^{2}-\frac{1}{3}\right).$ (58)
@@
$\begin{cases}\frac{d_{u}}{L^{2}}u_{2,1}^{\prime\prime}+u_{2,1}f^{\prime}\left(u_{2,0}\right)=\frac{2}{L}f\left(u_{2,0}\right)&0<x<1\\ u_{2,1}^{\prime}(1)=0&\\ u_{2,1}(0)=u_{1,1}(1)=\frac{1}{3L}u_{2,0}^{\prime}(0),\end{cases}$ (59)
@@
$\begin{cases}d_{u}u_{1,2}^{\prime\prime}=-f\left(\frac{\gamma}{\alpha}\right)+\frac{\beta\gamma}{\alpha}v_{0}+\frac{\alpha}{2L^{2}\gamma}\left(u_{2,0}^{\prime}(0)\right)^{2}\left(x^{2}-\frac{1}{3}\right)&0<x<1\\ u_{1,2}^{\prime}(0)=0&\\ u_{1,2}^{\prime}(1)=\frac{1}{L}u_{2,1}^{\prime}(0).\end{cases}$ (60)
@@
Using (54) and (62), and letting $w_{1}$ and $w_{2}$ be $u_{2,0}$ and $u_{2,1}$ represented in our original coordinates, the result follows. ∎
Depending on the parameters $\alpha$, $\beta$, $\gamma$, $f$, $d_{u}$, $d_{v}$, and $L$, the value of $V^{\prime}(a)$ may be positive or negative. This leads to the surprising result that the total predator population may actually be higher at $a=0$ than for small $a>0$. We will examine this phenomenon in detail in Section 6.
6. Further Properties and Open Problems
@@
We have explored several properties of solutions to the model (1) analytically. There are many further properties of solutions left to be studied. In this section, we demonstrate several such properties numerically and provide conjectures to motivate future analytical studies.
6.1. Computations of equilibria 1D
@@
Of principal interest is how solutions to (1) depend on $A$ and $\Omega$. Theorems 1.3 and 1.4 give conditions for existence of coexistence solutions, and in Section 5, we consider the case that asymptotic case when $A$ is small. But to explore the dependence of solutions on $A$ and $\Omega$ in generality, we turn to numerical solutions. So that this dependence may be understood in the simplest possible setting, we solve (1) numerically in 1D. Specifically, we consider the setting where $\Omega=(0,L)$ for some $L>0$ and $A=(0,a)$ for some $0<a<L$. Then the question of how solutions depend on $A$ and $\Omega$ reduces simply to the question of how they depend on the two numbers $a$ and $L$. However, even in this simple setting, we see that the model (1) gives rise to a remarkable variety of interesting solution behaviors. We show many such behaviors in figures in this section. Table 1 lists the parameters used to generate the solutions shown in each figure. The bistable growth function $f:[0,1]\to\mathbb{R}$ we use for these simulations is
@@
To begin, we examine the long-time behavior of solutions $v(t,x)$ and $u(t,x)$. We observe four types of long-time behavior.
  1. (1)
    Coexistence equilibrium. Solutions may approach a locally stable coexistence equilibrium $(\hat{u}(x),\hat{v}(x))$ as $t\to\infty$. Such an equilibrium solution is shown in Figure 2(a). Convergence to coexistence equilibria occurs in many parameter regimes, but seems always to occur if the length LaL-a of the exclusion zone is large enough and if the initial prey density is large enough to avoid immediate extinction.
  2. (2)
    Extinction. Predators may drive the prey to extinction, and then die off themselves. This occurs when the initial prey density is sufficiently small and the initial predator density is sufficiently large that, on a timescale short relative to the characteristic predator timescale $1/\gamma$, the predators drive the prey beneath their Allee threshold. This is difficult to achieve if the length LaL-a of the exclusion zone is large because the prey in the exclusion zone cannot be directly controlled by the predators. But as we shall see, when LaL-a is small, numeric solutions suggest that in many parameter regimes, the trivial solution is the only asymptotically stable trajectory.
  3. (3)
@@
  5. (5)
    Chaotic behavior. If the length $a$ of the predator zone is larger than the multiple waves of prey can be present simultaneously (but still such that LaL-a is large enough to allow the prey to persist), these waves may interfere leading to unpredictable and chaotic behavior. We hypothesize that as $a$ increases, the periodic solutions undergo a cascade of period-doubling bifurcations leading eventually to a chaotic attractor. A more detailed study of this phenomenon will be presented in forthcoming work.
@@
Figure 2. In some parameter regimes, solutions to (1) approach a coexistence equilibrium (left). In others, solutions approach a limit cycle (right).
Figure 3. These contour plots show the population densities of prey (top) and predators (bottom) depending on time and space in as solutions of (1) approach a limit cycle.
@@
$\overline{U}(a)$ $=\lim_{t\to\infty}\frac{1}{T}\int_{0}^{L}u(t,x)\,dx\quad\quad$ $\overline{V}(a)=\lim_{t\to\infty}\frac{1}{T}\int_{0}^{L}v(t,x)\,dx$ (67)
@@
  1. (1)
    Figure 4 show three features frequently seen in limiting profiles:
@@
    • Hopf bifurcation. There is a value $a_{\text{hopf}}$ such that for $a<a_{\text{hopf}}$, solutions approach coexistence equilibria while for $a>a_{\text{hopf}}$, solutions approach limit cycles, until $a$ becomes large enough for another bifurcation to occur.
    • Sudden extinction. At $a=a_{\text{ext}}$, a saddle-node bifurcation of limit cycles occurs. The stable limit cycle observed for $a_{\text{hopf}}<a<a_{\text{ext}}$ disappears (annihilating with another, unstable limit cycle), and solutions instead approach extinction.
  2. (2)
@@
    • Local maximum in the thin-limit. The limiting total predator approaches a local maximum in the thin-limit $a\to 0^{+}$. This was predicted in Section 5 due to the indeterminate sign of $v_{0}$ in (62). Numerically, we observe that the sign of $\lim_{a\to 0^{+}}\overline{V}(a)$ seems to match the sign of $\frac{\gamma}{\alpha}-\theta$, particularly when $L$ is large.
    • Global maximum at extinction. Remarkably, after reaching a local minimum, $\overline{V}(a)$ increases right up until the point of saddle-node bifurcation of limit cycles at $a_{\text{ext}}$. Thus, $a_{\text{max}}=a_{\text{ext}}$ (or, at least these two values are within the numerical step size used for $a$ in this computation).
  3. (3)
@@
    • Global maximum in the thin-limit. Not only does $\overline{V}(a)$ approach a local maximum as $a\to 0^{+}$, but this is actually the global maximum. This means that the best situation for the predators is if they cluster together, leaving the largest possible exclusion zone for the prey.
@@
Figure 4. A plot of the limiting profiles exhibiting an interior maximum, Hopf bifurcation, and extinction.
Figure 5. A plot of the limiting profiles exhibiting a local maximum in the thin-limit, a Hopf bifurcation and a global maximum at extinction.
Figure 6. A plot of the limiting profiles exhibiting a global maximum in the thin-limit and no Hopf bifurcation.
@@
Table 1. Parameter values used to generate each figure.
@@
  • In every case, $\overline{U}(a)$ is decreasing in $a$, indicating that increasing $a$ is always detrimental to the prey. On the other hand, the predators may do best in a small territory, a large territory, or an intermediate territory depending on the parameter values. It is expected that the situation in higher dimensions is even more complicated, with the limiting predator population depending not only on territory size but also shape. Even in one dimension, the question remains whether or not more predators can be supported within a single interval of length $a_{\text{max}}$, or several disjoint intervals whose lengths sum to $a_{\text{max}}$ with exclusion zones between them. These questions will be the topics of further research.
  • For small enough $a$, solutions always tend to converge to coexistence equilibria (assuming the initial prey density is large enough to avoid extinction), and the limiting total population does not vanish in the limit of small $a$:

### v2

@@ 4 Thin-limit analysis in 1D
on $(x_{0},L)$. But since $x_{0}\in(0,a_{0})$ is arbitrary, we conclude that $\hat{u}$ satisfies (35) on $(0,L)$.
From (34), we may also calculate
$\hat{u}^{\prime}(L)$ $=d-\frac{1}{d_{u}}\int_{x_{0}}^{L}f(\hat{u}(s))\,ds$
$=d-\lim_{n\to\infty}\frac{1}{d_{u}}\int_{x_{0}}^{L}f(u_{n}(s))\,ds$
$=d+\lim_{n\to\infty}\int_{x_{0}}^{L}u_{n}^{\prime\prime}\,ds$
$=d+\lim_{n\to\infty}-u_{n}^{\prime}(x_{0})$
$=0.$
Now we show that $\hat{u}(0)=\gamma/\alpha$. Let $2\delta=\hat{u}(0)-\gamma/\alpha$. If $\delta\neq 0$, then either $\delta>0$ or $\delta<0$. Suppose $\delta>0$. Letting $\Lambda$ be an upper bound for both $\|v_{a}\|_{L^{1}}$ and $\|u_{a}\|_{C^{1}}$, take $n$ sufficiently large that $a_{n}<\delta/\Lambda$. By the Mean Value Theorem, we must have $u_{n}>\frac{\gamma}{\alpha}+\delta$ for all $x\in(0,a_{n})$. Then, with $v_{n}:=v_{a_{n}}$,
$0=\int_{0}^{a_{n}}\left(v_{n}^{\prime\prime}+(\alpha u_{n}-\gamma)v_{n}\right)>\Lambda\alpha\delta>0,$
a contradiction. The $\delta<0$ case is exactly similar. We conclude that $\delta=0$ and thus $\hat{u}(0)=\frac{\gamma}{\alpha}$.
Thus, we have shown that $\hat{u}$ is a solution to (30). For each $n$, we know that
$u_{n}(L)>\zeta_{a_{n},L}(L)>\zeta_{a_{0},L}(L)>\zeta_{a_{0},L}(L_{0})>\zeta_{a_{0},L_{0}}(L_{0})>1-\eta_{0},$ (36)
from which we conclude that $1-\eta_{0}<\hat{u}(L)<1$. But due to Proposition 4.3, there is only one solution $u$ to (30) with $1-\eta_{0}<u(L)<1$, and we have assumed $\hat{u}\neq u$. This is a contradiction, and thus we conclude that $\lim_{a\to 0}u_{a}=u$ in $C^{0}$. We may repeat the derivation of (34) to see that $u_{n}$ converges to $u$ in $C^{2}(x_{0},L)$ for each $x_{0}\in(0,L)$. Finally, from (32), we have in the limit as $a\to 0$,
$\lim_{a\to 0}\int_{0}^{a}v_{a}\,dx=\frac{\alpha}{\beta\gamma}\int_{0}^{L}f(u)\,dx=-\frac{d_{u}\alpha}{\beta\gamma}\int_{0}^{L}u^{\prime\prime}\,dx=\frac{d_{u}\alpha}{\beta\gamma}u^{\prime}(0).$ (37)
Integrating (30) from 0 to $L$, we see that
$u^{\prime}(0)=\sqrt{\frac{2}{d_{u}}F(u(L)}.$
Since $u(L)>1-\eta_{0}>\theta^{\prime}$, each of $F(u(L))$, $u^{\prime}(0)$, and $\lim_{a\to 0}\int_{0}^{a}v_{a}\,dx$ are positive. ∎
From Theorem 4.4, we see that as $a\to 0^{+}$, the total predator population of coexistence equilibria for (1) with $\Omega=(0,L)$ (for large $L$) and $A=(0,a)$ approaches a finite, positive limit. This holds even if there is not a unique coexistence equilibrium $(u_{a},v_{a}$) for each $a$ sufficiently small, but if there is a sufficiently regular family of such equilibria, we could compute an asymptotic expansion for the total predator population:
$V(a)=\int_{0}^{a}v_{a}\,dx=V(0)+V^{\prime}(0)a+O(a^{2}).$
Theorem 4.4 already derives the constant term $V(0)$, so it remains to derive the linear coefficient $V^{\prime}(0)$. We will do this using formal asymptotic expansions, assuming sufficient regularity of the map $a\mapsto(u_{a},v_{a})$. In fact, this regularity is limited as $u_{a}$ is of necessity not twice differentiable at $x=a$. Instead, we will consider regularity of $u_{a}$ in $(0,a)$ and $(a,L)$ separately. To facilitate this, we perform a change of coordinates for solutions $u_{a}$ and $v_{a}$. Letting $u_{1}=u_{a}\big|_{(0,a)}$ and $u_{2}=u_{a}\big|_{(a,L)}$, we apply the coordinate transformation $x\to x/a$ to $u_{1}$ and $v_{a}$ and the coordinate transformation $x\to(x-a)/(L-a)$ to $u_{2}$ to obtain equations for $u_{1}$, $u_{2}$, and $v=v_{a}$ (we drop the subscript $a$ for convenience) in these new coordinates:
@@
We consider the following ansätze for the asymptotic expansions of $u_{1}$, $u_{2}$, and $v$ in small $a$, informed by Theorem 4.4:
$\begin{split}u_{1}(x)&=\frac{\gamma}{\alpha}+au_{1,1}(x)+a^{2}u_{1,2}(x)+O(a^{3})\\ u_{2}(x)&=w_{1}(Lx)+au_{2,1}(x)+O(a^{2})\\ v(x)&=\frac{1}{a}\frac{d_{u}\alpha}{\beta\gamma}w_{1}^{\prime}(0)+v_{0}(x)+av_{1}(x)+a^{2}v_{2}(x)+O(a^{3}),\end{split}$ (39)
Where $w_{1}$ is the solution to (30) guaranteed by Proposition 4.3. Note that the expansion for $v$ leads in order $a^{-1}$ since we expect the predator density to blow up as $a\to 0$ while the total predator population approaches a positive constant due to Theorem $4.$
Substituting (39) into (38) and comparing terms of like order in $a$, we see that $u_{1,1}$ solves
$\begin{cases}u_{1,1}^{\prime\prime}=w_{1}^{\prime}(0)&0<x<1\\ u_{1,1}^{\prime}(0)=0&\\ u_{1,1}^{\prime}(1)=w_{1}^{\prime}(0)\end{cases}$ (40)
We observe that the solvability condition for this Neumann problem is already exactly met. We can determine $u_{1,1}$ up to an additive constant:
$u_{1,1}(x)=\frac{1}{2}w_{1}^{\prime}(0)x^{2}+c.$ (41)
@@
$\begin{cases}d_{v}v_{2}^{\prime\prime}=-\frac{d_{u}\alpha^{2}}{\beta\gamma}w_{1}^{\prime}(0)u_{1,1}&0<x<1\\ v_{2}^{\prime}(0)=v_{2}^{\prime}(1)=0.\end{cases}$ (42)
We must have $\int_{0}^{1}v_{2}^{\prime\prime}\,dx=-\frac{d_{u}\alpha^{2}}{\beta\gamma}w_{1}^{\prime}(0)\int_{0}^{1}u_{1,1}\,dx=0$. Thus, we find
$u_{1,1}(x)=\frac{1}{2}w_{1}^{\prime}(0)\left(x^{2}-\frac{1}{3}\right).$ (43)
@@
$\begin{cases}\frac{d_{u}}{L^{2}}u_{2,1}^{\prime\prime}+u_{2,1}f^{\prime}\left(w_{1}(Lx)\right)=\frac{2}{L}f\left(w_{1}(Lx)\right)&0<x<1\\ u_{2,1}^{\prime}(1)=0&\\ u_{2,1}(0)=\frac{1}{3}w_{1}^{\prime}(0),\end{cases}$ (44)
@@
$\begin{cases}d_{u}u_{1,2}^{\prime\prime}=-f\left(\frac{\gamma}{\alpha}\right)+\frac{\beta\gamma}{\alpha}v_{0}+\frac{\alpha}{2L\gamma}\left(w_{1}^{\prime}(0)\right)^{2}\left(x^{2}-\frac{1}{3}\right)&0<x<1\\ u_{1,2}^{\prime}(0)=0&\\ u_{1,2}^{\prime}(1)=\frac{1}{L}u_{2,1}^{\prime}(0).\end{cases}$ (45)
@@
Using (47), and letting $w_{2}$ be $u_{2,1}$ represented in our original coordinates, we obtain
where $w_{1}$ is the solution to (30) guaranteed by Proposition 4.3 and $w_{2}$ is a solution to
$\begin{cases}d_{u}w_{2}^{\prime\prime}+w_{2}f^{\prime}(w_{1})=\frac{2}{L}f(w_{1})&0<x<L\\ w_{2}^{\prime}(L)=0&\\ w_{2}(0)=\frac{1}{3}w_{1}^{\prime}(0).\end{cases}$ (49)
Depending on the parameters $\alpha$, $\beta$, $\gamma$, $f$, $d_{u}$, $d_{v}$, and $L$, the value of $V^{\prime}(a)$ may be positive or negative. The following Theorem shows that when $L$ is large, the value of $V^{\prime}(a)$ approximately proportional to $f(\gamma/\alpha)$.
Theorem 4.5.
In the large $L$ limit, the linear coefficient in (48) is
$\lim_{L\to\infty}V^{\prime}(0)=\lim_{L\to\infty}\frac{\alpha}{\beta\gamma}\left(d_{u}w_{2}^{\prime}(0)+f\left(\frac{\gamma}{\alpha}\right)\right)=\frac{2}{3}\frac{\alpha}{\beta\gamma}f\left(\frac{\gamma}{\alpha}\right).$ (50)
Proof.
Let $w_{1}$ be the monotone solution to (30) in $[0,L]$ guaranteed by Proposition 4.3 for large $L$. Let $w_{2}$ be the solution to (49). We begin by showing that $w_{2}(L)$ vanishes as $L\to\infty$. We may differentiate $d_{u}w_{1}^{\prime\prime}+f(w_{1})=0$ to see that $w_{1}^{\prime}$ solves the linear equation
$d_{u}(w_{1}^{\prime})^{\prime\prime}+f^{\prime}(w_{1})w_{1}^{\prime}=0.$ (51)
Next, we will introduce two auxiliary functions. Define
$z(x)=w_{2}(x)+\left(\frac{x}{L}-\frac{1}{3}\right)w_{1}^{\prime}(x)$
Then $z$ solves
$d_{u}z^{\prime\prime}+f^{\prime}(w_{1})z$ $=d_{u}w_{2}^{\prime\prime}+d_{u}\left(\frac{x}{L}-\frac{1}{3}\right)w_{1}^{\prime\prime\prime}+\frac{2d_{u}}{L}w_{1}^{\prime\prime}+f^{\prime}(w_{1})w_{2}+\left(\frac{x}{L}-\frac{1}{3}\right)f^{\prime}(w_{1})w_{1}^{\prime}$
$=\left(d_{u}w_{2}^{\prime\prime}+f^{\prime}(w_{1})w_{2}\right)+\left(\frac{x}{L}-\frac{1}{3}\right)\left(d_{u}w_{1}^{\prime\prime\prime}+f^{\prime}(w_{1})w_{1}^{\prime}\right)-\frac{2}{L}f(w_{1})$
$=0.$
We also have $z(0)=w_{2}(0)-w_{1}^{\prime}(0)/3=0$ and
$z^{\prime}(L)=w_{2}^{\prime}(L)+\frac{1}{L}w_{1}^{\prime}(L)+\frac{2}{3}w_{1}^{\prime\prime}(L)=-\frac{2}{3}\frac{f(w_{1}(L))}{d_{u}}.$
Our next auxiliary function is $y$ given by
$y=\frac{\partial w_{1}}{\partial L}.$
Differentiating $d_{u}w_{1}^{\prime\prime}+f(w_{1})=0$ with respect to $L$, we see that that $y$ solves the same equation as $w_{1}^{\prime}$ and $z$: $d_{u}y^{\prime\prime}+f^{\prime}(w_{1})y=0$. For boundary conditions, we have $y(0)=0$ since $w_{1}(0)=\gamma/\alpha$ independent of $L$. We may differentiate $w_{1}^{\prime}(L)=0$ with respect to $L$ (using the chain rule) to obtain $y^{\prime}(L)+w_{1}^{\prime\prime}(L)=0$, from which we conclude
$y^{\prime}(L)=\frac{f(w(L))}{d_{u}}.$
Since $y$ and $z$ both satisfy the same linear differential equation and $y(0)=z(0)=0$, we conclude that they are scalar multiples of each other. From their boundary conditions at $L$, we see that $z=-\frac{2}{3}y$, and therefore
$w_{2}(L)=w_{2}(L)+\frac{2}{3}w_{1}^{\prime}(L)=z(L)=-\frac{2}{3}y(L).$ (52)
Thus, to show that $w_{2}(L)$ is bounded, it suffices to show that $y(L)$ is bounded. Denote $w_{1}(L)$ by $q$ as in the proof of Proposition 4.3 (in Appendix A). Since $q$ depends on $L$, we may write it as $q=q(L)$. Then we have $y(L)=q^{\prime}(L)$. In the proof of Proposition 4.3, we see that for large $L$, this function $q(L)$ is invertible and we may write $L=\mathcal{L}(q)$ as (66). Moreover, we also showed in this same proof that for sufficiently large $q$ (or equivalently, sufficiently large $L$), $\mathcal{L}$ is increasing and
$\lim_{q\to 1}\mathcal{L}(q)=\lim_{q\to 1}\mathcal{L}^{\prime}(q)=\infty.$
Therefore,
$\lim_{L\to\infty}-\frac{3}{2}w_{2}(L)=\lim_{L\to\infty}y(L)=\lim_{L\to\infty}q^{\prime}(L)=\frac{1}{\lim_{q\to 1}\mathcal{L}^{\prime}(q)}=0.$
Now, we derive the asymptotic value of $w_{2}^{\prime}(0)$ for large $L$. Differentiating $d_{u}w_{1}^{\prime\prime}+f(w_{1})=0$, we see that $w_{1}^{\prime}$ satisfies
$d_{u}w_{1}^{\prime\prime\prime}+f^{\prime}(w_{1})w_{1}^{\prime}=0.$ (53)
Multiplying (49) by $w_{1}^{\prime}$ and multiplying (53) by $w_{2}$, and subtracting the second from the first, we have
$d_{u}(w_{2}^{\prime\prime}w_{1}^{\prime}-w_{1}^{\prime\prime\prime}w_{2})=\frac{2}{L}f(w_{1})w_{1}^{\prime}$ (54)
Integrating by parts from 0 to $L$, we have
$d_{u}\left(w_{2}^{\prime}(L)w_{1}^{\prime}(L)-w_{2}^{\prime}(0)w_{1}^{\prime}(0)-w_{1}^{\prime\prime}(L)w_{2}(L)+w_{1}^{\prime\prime}(0)w_{2}(0)\right)=\frac{2}{L}(F(w_{1}(L))-F(\gamma/\alpha)).$ (55)
We may simplify the left hand side while (since $F$ is bounded), the right hand vanishes as $L\to\infty$:
$\lim_{L\to\infty}\left(-d_{u}w_{2}^{\prime}(0)w_{1}^{\prime}(0)+f(w_{1}(L))w_{2}(L)-\frac{1}{3}f\left(\gamma/\alpha\right)w_{1}^{\prime}(0)\right)=0.$ (56)
Since both $f(w_{1}(L))$ and $w_{2}(L)$ vanish as $L\to\infty$ and $w_{1}^{\prime}(0)$ is bounded away from 0 (as it approaches $\sqrt{2F(1)}$ as $L\to\infty$), we conclude that
$\lim_{L\to\infty}w_{2}^{\prime}(0)=-\frac{1}{3d_{u}}f\left(\frac{\gamma}{\alpha}\right).$
Substituting this into (48), the result is proved. ∎
The surprising result of Theorem 4.5 is that the total predator population may actually be higher at $a=0$ than for small $a>0$. This means that if the predators are very efficient ($\gamma/\alpha<\theta$), then they are better off restricting themselves to a small territory as $V^{\prime}(a)<0$. Indeed, if highly efficient predators occupied a small interval $(0,a)$ with $a>0$, then they would consume most of the prey immediately when it diffuses into the predation zone, leaving most of the interval $(0,a)$ devoid of prey, and therefore devoid of predators as well. Consequently, the main effect of occupying an interval of length $a$ as opposed to a single point at $x=0$ is to effectively decrease the size of the prey domain, limiting the amount of prey available and decreasing the total predator population at equilibrium. We will examine this phenomenon numerically in Section 5.
5 Further Properties and Open Problems
@@
We have explored several properties of solutions to the model (1) analytically. There are many further properties of solutions left to be studied. In this section, we demonstrate several such properties numerically and provide conjectures to motivate future analytical studies. The purpose of this section is not to provide a rigorous numerical analysis, but rather to illustrate and complement the analytical results of the previous sections. We use numerical simulations to (i) visualize coexistence equilibria, (ii) explore how the size of the exclusion zone influences solution behavior, and (iii) highlight dynamical phenomena that suggest directions for future analytical study.
5.1 Computations of equilibria 1D
@@
Of principal interest is how solutions to (1) depend on AA and $\Omega$. Theorems 4.2 and 1.2 give conditions for existence of coexistence solutions, and in Section 4, we consider the asymptotic case when AA is small. But to explore the dependence of solutions on AA and $\Omega$ in generality, we turn to numerical solutions. So that this dependence may be understood in the simplest possible setting, we solve (1) numerically in 1D. Specifically, we consider the setting where $\Omega=(0,L)$ for some $L>0$ and $A=(0,a)$ for some $0<a<L$. Then the question of how solutions depend on AA and $\Omega$ reduces simply to the question of how they depend on the two numbers $a$ and $L$. However, even in this simple setting, we see that the model (1) gives rise to a remarkable variety of interesting solution behaviors. We show many such behaviors in figures in this section. Table 1 lists the parameters used to generate the solutions shown in each figure. The bistable growth function $f:[0,1]\to\mathbb{R}$ we use for these simulations is
@@
For these numerical simulations, we solved the time-dependent reaction–diffusion system using a method-of-lines approach. Space was discretized by finite differences on two meshes—one uniform mesh on $[0,a]$, and another uniform mesh on $[a,L]$, allowing the grid to conform to the interface $x=a$. The prey equation was discretized on the full domain $[0,L]$ while the predator equation was discretized only on $[0,a]$. Homogeneous Neumann boundary conditions for both equations were incorporated directly into the discrete Laplacian, and the interface $x=a$ was handled using the corresponding nonuniform finite-difference stencil. This yielded a sparse system of ODEs, which were integrated using SciPy’s implicit Radau solver, facilitated by the analytic Jacobian of the discretized system. Simulations were initialized from spatially homogeneous data with prey density $u\equiv 1$ and predator density $v\equiv 0.5$ on the predator domain. Total prey and predator populations were computed by numerical quadrature of the discrete solutions. When computing the “terminal population”, the average of total population over the last quarter of the simulation window was taken. These computations are intended to illustrate the analytical results, visualize representative solution behavior, and explore qualitative phenomena suggested by the model. A link to a repository containing the code used for the simulations in this section is found in the Data Declaration section below.
To begin, we examine the long-time behavior of solutions $v(t,x)$ and $u(t,x)$. We observe five types of long-time behavior.
  1. 1.
    Coexistence equilibrium. Solutions may approach a locally stable coexistence equilibrium $(\hat{u}(x),\hat{v}(x))$ as $t\to\infty$. Such an equilibrium solution is shown in Figure 2(a). Convergence to coexistence equilibria occurs in many parameter regimes, but seems to occur consistently if the length LaL-a of the exclusion zone is large enough and if the initial prey density is large enough to avoid immediate extinction.
  2. 2.
    Extinction. Predators may drive the prey to extinction, and then die off themselves. This occurs when the initial prey density is sufficiently small and the initial predator density is sufficiently large that, on a timescale short relative to the characteristic predator timescale $1/\gamma$, the predators drive the prey beneath their Allee threshold. This is difficult to achieve if the length LaL-a of the exclusion zone is large because the prey in the exclusion zone cannot be directly controlled by the predators. But as we shall see, when LaL-a is small, numeric solutions tend to converge to the trivial equilibrium.
  3. 3.
@@
  5. 5.
    Chaotic behavior. If the length $a$ of the predator zone is larger than the multiple waves of prey can be present simultaneously (but still such that LaL-a is large enough to allow the prey to persist), these waves may interfere leading to apparently irregular (and possibly chaotic) behavior. We hypothesize that as $a$ increases, the periodic solutions undergo a cascade of period-doubling bifurcations leading eventually to a chaotic attractor. A more detailed study of this phenomenon will be presented in forthcoming work.
@@
Figure 2: In some parameter regimes, solutions to (1) approach a coexistence equilibrium (left). In others, solutions approach a limit cycle (right).
Figure 3: These contour plots show the population densities of prey (top) and predators (bottom) depending on time and space as solutions of (1) approach a limit cycle.
@@
$\overline{U}(a)$ $=\lim_{t\to\infty}\frac{1}{T}\int_{0}^{L}u(t,x)\,dx\quad\quad$ $\overline{V}(a)=\lim_{t\to\infty}\frac{1}{T}\int_{0}^{a}v(t,x)\,dx$ (61)
@@
  1. 1.
    Figure 4 shows three features frequently seen in limiting profiles:
@@
    • Hopf bifurcation. There is a value $a_{\text{hopf}}$ such that for $a<a_{\text{hopf}}$, solutions appear to approach coexistence equilibria while for $a>a_{\text{hopf}}$, solutions appear to approach limit cycles, until $a$ becomes large enough for another bifurcation to occur. This suggests that a Hopf bifurcation is present.
    • Sudden extinction. At $a=a_{\text{ext}}$, the periodic behavior suddenly vanishes and solutions instead approach extinction (possibly a saddle-node bifurcation of limit cycles occurs here).
  2. 2.
@@
    • Local maximum in the thin-limit. The limiting total predator approaches a local maximum in the thin-limit $a\to 0^{+}$. This was predicted in Section 4 due to the indeterminate sign of $v_{0}$ in (47). This reflects the result of Theorem 4.5, which predicts that the predator population is locally maximized in the limit $a\to 0$ if $\gamma/\alpha<\theta$.
    • Global maximum at extinction. Remarkably, after reaching a local minimum, $\overline{V}(a)$ increases right up until the point of saddle-node bifurcation of limit cycles at $a_{\text{ext}}$. Thus, $a_{\text{max}}$ appears to coincide with $a_{\text{ext}}$ (or, at least these two values are within the numerical step size used for $a$ in this computation).
  3. 3.
@@
    • Global maximum in the thin-limit. Not only does $\overline{V}(a)$ approach a local maximum as $a\to 0^{+}$, but this is actually the global maximum. This suggests that the best situation for the predators is if they cluster together, leaving the largest possible exclusion zone for the prey.
@@
Figure 4: A plot of the limiting profiles exhibiting an interior maximum, Hopf bifurcation, and extinction.
Figure 5: A plot of the limiting profiles exhibiting a local maximum in the thin-limit, a Hopf bifurcation and a global maximum at extinction.
Figure 6: A plot of the limiting profiles exhibiting a global maximum in the thin-limit and no Hopf bifurcation.
@@
Table 1: Parameter values used to generate each figure.
@@
  • In every case examined, $\overline{U}(a)$ is decreasing in $a$, indicating that increasing $a$ is always detrimental to the prey. On the other hand, the predators may do best in a small territory, a large territory, or an intermediate territory depending on the parameter values. It is expected that the situation in higher dimensions is even more complicated, with the limiting predator population depending not only on territory size but also shape. Even in one dimension, the question remains whether or not more predators can be supported within a single interval of length $a_{\text{max}}$, or several disjoint intervals whose lengths sum to $a_{\text{max}}$ with exclusion zones between them. These questions will be the topics of further research.
  • For small enough $a$, solutions are consistently observed to converge to coexistence equilibria (assuming the initial prey density is large enough to avoid extinction), and the limiting total population does not vanish in the limit of small $a$: